OpenStudy (anonymous):

A 1,200-watt water heater is plugged into a 120-volt outlet and used for 1.5 hours. How much current runs through the vacuum? @KonradZuse

OpenStudy (konradzuse):

Power is P(t) * V(t)

OpenStudy (anonymous):

all i can find online are calculators to answer the problem. i cant find a formula so i can learn how to do this :(

OpenStudy (konradzuse):

I(t)*

OpenStudy (konradzuse):

so 1200 = i(1.5) * 120(1.5)

OpenStudy (konradzuse):

1200/1.5 = i * 120

OpenStudy (anonymous):

so 1200/ 1.5 is 800______ 800=i*120_________ \[\frac{ 800 }{ 120 }\] =i so i =6.7?

OpenStudy (anonymous):

thats not one of my answer choices though

OpenStudy (anonymous):

is it 10?

OpenStudy (anonymous):

well, thanks anyway

OpenStudy (konradzuse):

didn't see your replies here, sorry. What are your choices?

OpenStudy (konradzuse):

oh wait.... 1.5 HOURS...........

OpenStudy (konradzuse):

so we need to do 60 seconds * 60 mins * 1.5 hours